### v2

# Knolling Bot: Learning Robotic Object Arrangement from Tidy Demonstrations

Abstract
Addressing the challenge of organizing scattered items in domestic spaces is complicated by the diversity and subjective nature of tidiness. Just as the complexity of human language allows for multiple expressions of the same idea, household tidiness preferences and organizational patterns vary widely, so presetting object locations would limit the adaptability to new objects and environments. Inspired by advancements in natural language processing (NLP), this paper introduces a self-supervised learning framework that allows robots to understand and replicate the concept of tidiness from demonstrations of well-organized layouts, akin to using conversational datasets to train Large Language Models(LLM). We leverage a transformer neural network to predict the placement of subsequent objects. We demonstrate a “knolling” system with a robotic arm and an RGB camera to organize items of varying sizes and quantities on a table. Our method not only trains a generalizable concept of tidiness, enabling the model to provide diverse solutions and adapt to different numbers of objects, but it can also incorporate human preferences to generate customized tidy tables without explicit target positions for each object.
I INTRODUCTION
Designing a robot for household tasks has always presented unique challenges [1, 2, 3, 4]. Unlike industrial settings characterized by uniformity and limited object variety, household environments are filled with distinct objects. Recent progress in object rearrangement has shown robots can organize objects in partially arranged scenes, cluttered tabletops, and constrained environments[5, 6, 7]. While they provided effective approaches to solving task-specific problems, it is crucial to equip the robot with a task-agnostic planner to organize objects based on a generalized sense of tidiness.
The objects in our daily environments, such as desks and houses, vary constantly. In such dynamic settings, providing specific goals for each object is impractical and limits the robot’s generalizability. Instead, robots should be capable of organizing the environment without relying on specific target positions or constant human supervision. This requires the robot to understand and apply a broad concept of organization that extends beyond the specifics of any single task. The necessary adaptability is similar to human-like cognition, where decisions are made not only based on spatial arrangement but also considering the aesthetic and functional aspects of tidiness.
“Knolling” is one such concept intrinsic to humans, referring to the intuitive ability to organize items in a manner that is both aesthetically pleasing and space-efficient (Fig. 1A). Our solution lies in decoupling the cognitive model – which encapsulates the essence of knolling – from the other modules, such as the visual perception system and motor controllers[8]. This division enhances the modularity and interpretability of the entire system compared to only training a single policy mapping from observation to action.
Figure 1: Objects for real-world experiments in a knolling layout. A) A batch of small items, including daily necessities. B) A batch of big items fabricated by a 3d printer.
Figure 2: A) Challenges in rearrangement tasks with multiple solutions. From left to right: the initial state of the work area with an unplaced yellow motor. Three proposed placement options (1, 2, 3) for the yellow motor, considering factors such as object category, color similarity, and spatial efficiency. Placement by a regression-based model optimizing for minimal loss across three solutions, leading to undesirable results: placing the motor on a utility knife. B) Knolling pipeline. The left side of the figure displays a cluttered desktop with various objects such as batteries, erasers, electronic components, and other daily necessities in the lab. Our robot initiates the knolling process after detecting and identifying these objects through a camera. The right side of the figure depicts the outcome of this process, presenting a tidy, well-organized desktop. This transformation exemplifies the robot’s ability to apply the knolling model, execute a tidying task, and create a pleasing and space-efficient arrangement.
Given the subjectivity of tidiness and the multiple optimal solutions that could satisfy different individuals’ preferences, knolling is an abstract problem without any definitive standards. Consider the scenario depicted in Fig.2A, where a robotic arm performs a task to integrate a yellow motor into a partially organized environment. The decision on where to place the motor can be depended on various factors, including object category, color, and the need for space efficiency. Training a model solely on regression to predict target positions based on object states might lead to local optimal outcomes. For instance, a model optimized to minimize the loss across different placement preferences could converge to an average of the three potential target positions, resulting in the motor being placed overlapped with a utility knife. This is just an example of a one-step rearrangement of a partially organized scenario. A wrong-placed object may cause the next step of rearrangement to become worse because the errors have accumulated. Moreover, quantifying tidiness with a singular metric is challenging, as no simple equation can encapsulate the diverse aspects of what constitutes a “tidy” space.
In this work, we propose a self-supervised learning framework for modeling tidiness through demonstrations, akin to training a Large Language Model (LLM) with human conversational datasets[9, 10, 11, 12]. We draw parallels between knolling objects and language, where objects act as individual “words” that can be combined in various ways to form “sentences” with identical meanings. Besides, transformer architecture can handle varying input and output sizes with autoregression, which fits the knolling task as the number of objects in the environment is arbitrary[13, 14]. Leveraging the Gaussian Mixture Model (GMM), we address the challenges of multi-label prediction to perform the knolling task[15].
This research aims to contribute to bridging this gap by imbuing a robot with an understanding of the human concept of neatness and developing an integrated pipeline to demonstrate how to deploy it to perform real-world knolling tasks. The pipeline shown in Fig.2B comprises three stages: a knolling model, a visual perception model, and a robot arm controller. The knolling model, based on a transformer architecture, predicts the target positions of objects. The visual perception model, based on our customized YOLO v8, detects objects from an RGB image [16]. Finally, the robot arm controller, leveraging the outputs from the previous stages, guides the robot arm to execute the knolling tasks.
The key contributions of this work are as follows:
1. We propose a novel self-supervised learning framework for modeling the concept of tidiness. Inspired by advancements in natural language processing, our approach treats objects as “words” that can be organized into aesthetically pleasing and space-efficient “sentences”, allowing the model to learn generalizable tidiness patterns from demonstrations.
2. We leverage transformer architectures and Gaussian Mixture Models to address the multi-label prediction challenges inherent in knolling tasks, where multiple optimal solutions may exist. This enables our model to generate diverse, tidy arrangements tailored to different preferences.
3. We present an integrated end-to-end pipeline that combines the learned knolling model with visual perception and robotic control modules. This allows for deploying knolling capabilities on a real robotic system to organize cluttered environments with varying object quantities and types.
4. Through our experiments, we demonstrate the effectiveness of our approach in generating tidy object layouts, its generalization to unseen objects, and its ability to adapt to user preferences, highlighting the potential for deploying such systems in dynamic household environments.
II RELATED WORK
The domain of robot learning and manipulation tasks has seen significant advancements in recent years. Various manipulation tasks, such as wiping and polishing, stacking, peg-in-hole, and pick-and-place tasks, have been explored[17, 18, 19, 20, 21]. In the realm of stacking tasks, Lee et al. used offline reinforcement learning (RL) to improve upon existing policies for robotic stacking of objects with complex geometry, while Furrer et al. proposed an algorithm for suggesting stable poses for stacking, validated through a real-world autonomous stacking workflow[22, 23]. For pick-and-place tasks, Gualtieri et al. proposed a deep reinforcement learning approach, and Zeng et al. introduced the Transporter Network for vision-based manipulation tasks[24, 25].
In the domain of machine learning models, transformers have revolutionized many fields with their attention-based architecture[26, 27, 28, 29, 30, 31]. The success of transformer models in NLP has prompted exploration in their application to robotic tasks[32, 33, 34]. Jangir et al. proposed the use of transformers with a cross-view attention mechanism for effective fusion of visual information from both cameras for RL policies[35]. Zhu et al. introduced VIOLA, which used a transformer-based policy to improve the robustness of deep imitation learning algorithms[36]. Shridhar et al. proposed PerAct, which uses the preceiver transformer to encode language goals and RGB-D voxel observations[37]. Jain et al. proposed the Transformer Task Planner, which could be pre-trained on multiple preferences, and Dasari et al. explored one-shot visual imitation learning using the Transformer architecture[38].
In the object rearrangement field, most recent works focused on rearranging objects based on explicit instructions[39, 40]. Some other works use a more general method to train a representation model or reward function to supervise the object rearrangement policy without explicit human supervision[41, 42]. LLM progresses, many works leverage the LLM to process natural language instructions from humans and provide robot goal locations for object rearrangement[43, 44, 45]. Instead of giving natural language instructions to the robot, some other works train a robot a command sense of how to place the objects through demonstrations[38, 46, 47]. Combining both approaches, Housekeep[48] provides a method that leverages LLM to train a common sense of tidiness for robots. While these works have contributed significantly to robot manipulation, we aim to provide a common sense of tidiness for robots through demonstrations. We are focusing on knolling tasks that rearrange messy objects on a desktop into a neat layout in the real world. Our approach allows for the handling of varying object types, sizes, and quantities, thus enabling the robot to generate an aesthetically pleasing and space-efficient arrangement of items, similar to human performance.
Figure 3: Knolling model learning framework: The first step in our pipeline involves recognizing and identifying the objects present on a table using the visual perception model. This model takes the input image and outputs n objects’ size (width and length) and position presented as a list. The Knolling Model utilizes a transformer-based model that takes the high-dimensional object state data produced by Positional Encoding as input. The model predicts N target positions after N iterative process.
III METHOD
Our approach of leveraging self-supervised learning on demonstrations of tidy arrangements to train a generalized knolling model is not only effective for desktop object organization but also scalable to broader tidiness tasks. Once a robot can learn the knolling task from tidy desktop demonstrations, the same approach could potentially be extended to larger environments. For example, well-organized room demonstrations could be used to train robots to perform general housekeeping and cleaning tasks. This is analogous to how large language models learn general language understanding from vast datasets of natural conversations and texts.
III-A Data Representation and Generation
We represent objects solely based on their dimensional parameters (width and length), as these are the fundamental properties governing spatial organization. Semantic attributes like color and category are non-differentiable and may introduce subjective biases, so we exclude them from the input data.
To generate a diverse dataset of tidy arrangements, we employ an optimization strategy that iteratively adjusts object placements to minimize the occupied area on the table, subject to different preferences (e.g., color, dimension, category). This stochastic process yields 2.4 million demonstrations spanning a wide variety of tidiness concepts and aesthetic preferences.
For objects with irregular shapes (Fig. 1B), we train a visual perception system to perform segmentation and compute the minimum bounding box, enabling our framework to accommodate unconventional geometries.
III-B Gaussian Mixture Model for Multi-target Learning
To address the multi-target nature of knolling tasks, where multiple valid arrangements are possible for a given set of objects as shown in Fig.2A, we employ Mixture Density Networks (MDNs) to model the potential distribution of object placements. MDNs combine a neural network with a Gaussian Mixture Model (GMM), allowing the network to predict not just a single target position, but a weighted mixture of Gaussian distributions representing the diverse set of plausible positions for each object.
Formally, the MDN component of our model predicts the parameters ${(\pi_{j},\mu_{j},\sigma_{j})}_{j=1}^{J}$ for a GMM with $J$ components, where $\pi_{j}$ are the mixing coefficients, $\mu_{j}$ are the means, and $\sigma_{j}$ are the standard deviations of the $J$ Gaussian distributions. For each object $i$, the predicted probability density function (PDF) over its possible placement positions $S_{i}$ is given by:
$p(S_{i})=\sum_{j=1}^{J}\pi_{ij}\mathcal{N}(S_{i}|\mu_{ij},\sigma_{ij}^{2})$ (1)
This formulation allows the model to capture the multiple knolling solutions, where identical object configurations can lead to different but equally valid tidy arrangements. During inference, we can sample from the predicted GMM to obtain diverse knolling predictions, or select the component with the highest mixing coefficient for a single deterministic prediction. Incorporating MDNs into our transformer architecture prevents the model from converging to averaging solutions and captures the inherent uncertainty and variability in tidiness preferences.
III-C Training Knolling Model
We used transformer architecture with encoder-decoder and autoregressive mechanism to iteratively predict object positions based on the object shape. The encoder learns all the shapes of the objects on the table, and the decoder is used to predict the target distributions. The encoder takes a list of the object’s width and length as input $O=\{o_{n}|o_{n}=(w_{n},l_{n}),n\in\mathbb{N}^{+}\}$. The knolling pipeline we presented is shown in Fig. 3.
Leveraging the autoregression of the transformer model, we initialize the model with fully-masked data ($M_{1},...,M_{n}$ in the Fig.3), which is then combined with the encoder’s output to predict the target position of the first object. Once the first object’s position is predicted, it replaces the first element of the masked data, and the model proceeds to predict the target position for the second object. This iterative process continues until all object positions are predicted. The knolling model can be expressed as:
$P_{n}=Dnc\left(Enc(O),\{P_{1},P_{2},\ldots,P_{n-1}\}\right),n\in\mathbb{N}^{+}$
We utilize 2 million data for training and 400k data for testing. The training process leverages a combination of five loss functions to train a model to predict object positions without overlaps while also promoting diversity in predictions:
1. Log-likelihood Loss ($L_{\text{ll}}$)
The Log-likelihood loss is a standard approach when working with GMM, which allows the model to express uncertainty and multi-targets in the predictions by assigning probabilities to a range of possible positions for each object. This is particularly useful for handling the diversity and complexity of knolling tasks, where the spatial arrangement of objects can vary significantly. The log-likelihood loss for each object $i$, considering a Gaussian Mixture Model (GMM) with $J$ components, is given by:
$L_{\text{ll}}=-\sum_{i=1}^{N}\log\left(\sum_{j=1}^{J}p_{ij}\cdot\mathcal{N}(S_{i}|\mu_{ij},\sigma_{ij}^{2})\right)$ (2)
where $N$ is the total number of objects, $p_{ij}$ is the weight of the $j$-th component for the $i$-th object, $\mathcal{N}(S_{i}|\mu_{ij},\sigma_{ij}^{2})$ is the probability density of the $i$-th object’s target position $S_{i}$ under the $j$-th Gaussian component with mean $\mu_{ij}$ and variance $\sigma_{ij}^{2}$.
2. MSE Loss ($L_{\text{MSE}}$)
The Mean Squared Error (MSE) Loss is an intuitive choice. It computes the squared difference between the sampled predicted position ($\hat{S}$) and the target position ($S$):
$L_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{S}_{i}-S_{i})^{2}$ (3)
where $\hat{S}_{i}$ is the sampled position from the predicted distribution for the $i$-th object, and $S_{i}$ is the actual target position. In a multi-target task, relying solely on MSE could lead the model to predict average positions in the presence of ambiguity, which is not desirable. The MSE Min Loss addresses this by allowing the model to choose the prediction (from among multiple Gaussian components) that is closest to the actual target position.
3. MSE Min Loss ($L_{\text{MSE Min}}$)
This encourages the model not only to predict accurately but also to account for the multiple solutions in the dataset. Considering $J$ components in the Gaussian mixtures for each prediction step, the MSE Min Loss is calculated as follows:
$L_{\text{MSE Min}}=\frac{1}{N}\sum_{i=1}^{N}\min_{j=1}^{J}\left((\hat{S}_{ij}-S_{i})^{2}\right)$ (4)
4. Overlapping Loss ($L_{\text{overlap}}$)
The Overlapping Loss penalizes predictions where objects overlap or cross boundaries. It acts as a form of regularization but with a very specific purpose: to ensure that the predicted positions do not result in physically impossible arrangements, such as overlapping objects or objects placed outside the designated workspace. This loss component is essential for maintaining the feasibility of the predicted knolling arrangements, ensuring that the model’s predictions adhere to spatial constraints inherent in the real-world application of knolling. The loss function can be expressed as:
$L_{\text{overlap}}=\sum_{i\neq k}\text{Penalty}(\hat{S}_{i},\hat{S}_{k},w_{n},l_{n},W,L)$ (5)
The penalty function calculates the overlap and out-of-boundaries areas based on their predicted positions ($\hat{S}_{i},\hat{S}_{j}$), the given dimensions of objects ($w_{n},l_{n}$) and workspace ($W,L$).
5. Entropy Loss ($L_{\text{entropy}}$)
This loss function encourages diversity in the predictions by measuring the entropy across the probabilities of the five Gaussian distributions for each prediction. The entropy loss is defined as:
$L_{\text{entropy}}=-\sum_{i=1}^{N}\sum_{j=1}^{J}p_{ij}\log(p_{ij})$ (6)
Figure 4: Examples of knolling messy tables with different numbers of objects. The figure shows ten examples of tables before and after the knolling process in the simulation.
6. Weights of Loss Components
We assigned each loss component a weight to balance the contributions of each loss during the training process. We used MSE Min loss and overlapping loss to evaluate the knolling performance.
III-D Knolling Pipeline
Our knolling pipeline comprises three modules: a knolling model, a visual perception model, and the robot arm controller for executing the pick-and-place task.
The visual perception model is a customized YOLO v8 model [16], fine-tuned to predict the information of the objects, including the segmentation, color, and category. We post-process the segmentation to a rectangular box and then extract the object state, including the center position, orientation, and dimensions. Detailed evaluation results and confusion matrices for the visual perception model in both simulated and real-world environments are provided in the supplementary materials. The data required for training our models is generated within the Pybullet simulation environment for pre-training and collected in the real world for fine-tuning. To prepare our Visual Perception Model for real-world deployment and address the simulation-to-reality gap, we apply the visual domain randomization technique to our data collection process [49]. This technique helps in enhancing the model’s adaptability and robustness by introducing variations and uncertainties that mimic real-world conditions. We manipulate several factors within the simulation environment, such as brightness and ground textures. Additionally, we also vary the appearance of the objects with different colors. Using this approach, we train our model to disregard extraneous noise and focus on essential features.
The robotic arm controller operates in four modes: movement, grasping/releasing, table sweeping, and object separation. It uses the predicted object positions from the knolling model and the current poses from the perception module to execute smooth pick-and-place operations, clearing occupied spaces through sweeping motions when necessary.
This integrated pipeline demonstrates how our self-supervised knolling model can be leveraged for practical robotic organization tasks, enabling autonomous tidying of cluttered environments without explicit human supervision or target specifications.
IV Experiments
To evaluate the performance and practical applicability of our proposed knolling framework, we conducted experiments in both simulated and real-world environments. Additionally, we performed quantitative analyses and ablation studies to validate the effectiveness of our approach and elucidate the contributions of different components.
IV-A Qualitative Evaluations in the Simulation
IV-A1 Handling Varying Object Quantities
In real-world scenarios, the number of objects on a surface can vary significantly. To assess our model’s ability to handle this variability, we conducted experiments with diverse object counts ranging from 2 to 10 items. Our model, leveraging its autoregressive nature, successfully generated tidy arrangements across all object quantities, as shown in Fig. 4. This adaptability to varying input sizes is a crucial aspect of our approach, enabling it to generalize to different scenarios without explicit retraining.
IV-A2 Generating Diverse Solutions Based on Preferences
Tidiness preferences can differ significantly among individuals, leading to multiple valid solutions for the same set of objects. To capture this diversity, our model decouples preferences from the input data (which consists solely of object dimensions) and instead encodes preferences through the order of the input sequence. By altering the input order based on criteria such as object category, color, or size, our model can generate distinct tidy arrangements tailored to specific preferences, as illustrated in Fig. 5. This flexibility to incorporate user preferences without modifying the model architecture is a key strength of our approach.
Figure 5: Knolling with preferences. This figure shows distinct layouts achieved by controlling the input order based on different criteria: 1. Categories: Place the same functionalities object together, 2. Shape: Big area priority, 3. Color: Group the objects that have the same colors.
Figure 6: A) Box knolling in the real world. In each test, we show four columns. Column 1: The initial state of the objects on a table, as captured by the overhead camera. Column 2: The same scenario as Column 1, with added key points and contour outlines indicating the detected objects. Column 3: Action snapshot of the robot executing the knolling task. Column 4: The final state of the workspace post-knolling, presenting an tidy table. B) Real-world Knolling Process with Different Object Numbers. This figure exhibits the practical application of our knolling model in four diverse scenarios. Each column corresponds to a different setup with a distinct number of objects (6, 8, 10, and 10). We show the initial messy state captured by the overhead camera and the organized layout after the knolling task is completed by the robot arm. These comparative visuals underline our robot’s proficiency in performing real-world knolling tasks across varied object quantities. C) For the same objects on the table, our robot can perform knolling tasks with different solutions based on preferences based on category, color, or shape.
IV-B Real-World Knolling Experiments
To demonstrate the practical applicability of our framework, we deployed our trained knolling model on a 5-DoF robotic arm (WidowX 200) equipped with an Intel Realsense D435 camera for visual perception. We randomly placed 6-10 boxes of varying sizes and colors within the robot’s workspace, simulating cluttered tabletop scenarios.
The integrated pipeline (Fig. 2) first captures an overhead image of the cluttered scene, which is processed by our visual perception module to detect objects, perform segmentation, and extract their dimensions and poses. The knolling model then predicts the target positions for each object based on their dimensions. Finally, the robotic arm controller executes smooth pick-and-place operations, guided by the predicted target positions and current object poses, to realize the tidy arrangements.
Fig. 6A showcases real-world knolling tasks performed on different box configurations, demonstrating our system’s capability to organize diverse object sets. Fig. 6B further illustrates the system’s adaptability to varying object quantities, successfully knolling scenes with 6, 8, and 10 objects. Additionally, Fig. 6C highlights the generation of distinct tidy arrangements for the same set of objects, reflecting different user preferences based on category, color, or shape. We encourage readers to view our supplementary video for a comprehensive visualization of these real-world knolling experiments.
IV-C Quantitative Evaluation and Ablation Study
To quantitatively assess the performance of our knolling model, we employed the L1 distance metric, which measures the absolute difference between the predicted and ground truth object positions. For deterministic results during evaluation, we selected the predicted Gaussian distribution based on minimum loss instead of the mixture weights. Our evaluation was conducted on a diverse test dataset comprising 20,000 samples for each object count ranging from 2 to 10 items, ensuring a comprehensive assessment across various knolling scenarios.
TABLE I: Comparisons between our methods and baselines
N objs. 2 4 6 8 10
MEAN 3.38E-04 2.40E-04 1.72E-04 2.08E-04 3.06E-04
OM STD 2.03E-04 2.71E-04 2.94E-04 4.13E-04 5.51E-04
MIN 8.52E-05 2.60E-06 3.40E-07 3.88E-06 1.36E-05
MAX 1.74E-03 2.16E-03 3.83E-03 4.95E-03 7.77E-03
MEAN 1.72E-02 2.09E-02 2.35E-02 2.62E-02 2.98E-02
LSTM STD 3.78E-03 3.53E-03 4.15E-03 3.88E-03 6.36E-03
MIN 8.37E-03 1.20E-02 1.43E-02 1.63E-02 1.86E-02
MAX 2.65E-02 3.10E-02 3.47E-02 4.37E-02 4.99E-02
MEAN 2.17E-01 1.45E-02 1.90E-01 2.55E-01 3.26E-01
MLP STD 1.91E-03 4.68E-03 2.42E-03 7.97E-03 8.95E-03
MIN 2.14E-01 6.98E-03 1.81E-01 2.38E-01 3.07E-01
MAX 2.23E-01 3.07E-02 2.00E-01 2.83E-01 3.57E-01
TABLE II: Ablation study
Model Test Performace Log-likelihood Loss MSE Loss MSE Min Loss Overlapping Loss Entropy Loss
OM 5.78E-04 $\pm$ 6.64E-04 -6.37E+00 $\pm$ 1.25E+00 4.91E-04 $\pm$ 6.25E-04 8.76E-03 $\pm$ 1.60E-02 9.23E-01 $\pm$ 9.55E-03 2.55E-02 $\pm$ 1.43E-02
OM w/o LL. Loss 7.24E-04 $\pm$ 7.53E-04 1.13E+02 $\pm$ 1.61E+02 6.34E-04 $\pm$ 7.05E-04 9.07E-03 $\pm$ 1.43E-02 9.01E-01 $\pm$ 2.25E-08 3.56E-02 $\pm$ 8.82E-03
OM w/o Pos. Loss 1.23E-02 $\pm$ 2.74E-03 -1.50E+00 $\pm$ 4.53E-01 1.05E-02 $\pm$ 2.47E-03 1.80E-01 $\pm$ 1.25E-01 1.06E+00 $\pm$ 4.09E-03 1.71E-01 $\pm$ 1.20E-02
OM w/o Pos. Min Loss 3.93E-02 $\pm$ 7.48E-03 3.51E-01 $\pm$ 5.64E-01 3.90E-02 $\pm$ 7.48E-03 2.32E-02 $\pm$ 3.37E-02 1.09E+00 $\pm$ 4.73E-03 1.14E-01 $\pm$ 1.13E-02
OM w/o Over. Loss 5.08E-02 $\pm$ 1.22E-02 -6.92E+00 $\pm$ 1.55E+00 3.40E-04 $\pm$ 4.07E-04 5.05E+00 $\pm$ 1.22E+00 9.66E-01 $\pm$ 2.44E-02 1.95E-02 $\pm$ 1.24E-02
OM w/o Ent. Loss 1.03E-03 $\pm$ 1.00E-03 -4.71E+00 $\pm$ 1.04E+00 9.07E-04 $\pm$ 9.43E-04 1.20E-02 $\pm$ 1.97E-02 1.08E+00 $\pm$ 5.70E-02 2.54E-02 $\pm$ 1.27E-02
Two baseline models were established to enable a relative performance comparison to Our Method (OM). The first baseline model utilized a Multilayer Perceptron (MLP) architecture, and the second baseline was designed based on the Long Short-Term Memory (LSTM) model. We use LSTM as a baseline because knolling can also be viewed as a sequence prediction problem where the arrangement of objects depends on their sequence. LSTMs have been extensively used for various sequence prediction tasks and have shown good performance. Thus, LSTMs serve as a more challenging baseline to benchmark the performance of our transformer-based model. Moreover, comparing our model with LSTM helps to highlight the benefits of using self-attention and auto-regression in the transformer for handling variable input and output sizes and multi-label problems. Just as we did with our model, we measured the L1 distance between the actual and predicted positions for each baseline model, with the results detailed in Table I. As a commitment to a fair evaluation, we have ensured that each model utilizes a similar amount of parameters: our model incorporates 87,458 parameters, the LSTM baseline uses 86,858 parameters, and the MLP baseline operates with 87,788 parameters. Our knolling model consistently outperforms the MLP and LSTM baselines in terms of the L1 error. This superiority of our model is evident in all parameters: mean L1 error, standard deviation, and minimum and maximum error. These results validate the effectiveness of our transformer-based model, which employs self-attention and auto-regression in handling varying input sizes and multi-label problems in knolling tasks.
Furthermore, we conducted an ablation study to evaluate the contributions of the individual loss components employed during training (Table II). The ablation results demonstrated the importance of each loss term, with their removal leading to performance degradation, justifying the need for a comprehensive loss function to capture the multifaceted nature of knolling tasks effectively. Additional experiments evaluating the model’s performance across varying dataset sizes and the effectiveness of the pre-training process are included in the supplementary material.
V CONCLUSIONS
In this work, we introduced a self-supervised learning framework for enabling robots to understand and replicate the abstract concept of tidiness, or “knolling,” from demonstrations of well-organized object arrangements.
Our approach leverages transformer architectures and Mixture Density Networks to model the inherent multi-modality of knolling tasks, where identical object configurations can lead to diverse but tidy arrangements. By training on a large dataset of knolling demonstrations without explicit human supervision or target positions, our model learns a generalizable concept of tidiness that can adapt to varying object quantities and incorporate user preferences.
Through experiments in both simulated and real-world environments, we demonstrated the effectiveness of our framework in generating aesthetically pleasing and space-efficient object arrangements. Our quantitative evaluations showed that our transformer-based model consistently outperformed baseline architectures, highlighting the advantages of self-attention and autoregressive mechanisms for handling variable input sizes and multi-label predictions.
Moreover, we presented an integrated pipeline that combines our learned knolling model with visual perception and robotic control modules, enabling the deployment of knolling capabilities on a real robotic system. Our real-world experiments showcased the system’s ability to organize cluttered environments with diverse object types and quantities, adapting to different user preferences based on criteria such as object category, color, or shape.
Future work could expand the training data to include demonstrations of well-organized rooms or environments, enabling robots to learn and apply overarching concepts of tidiness at a larger scale and facilitating autonomous housekeeping in dynamic household settings.